One of two complementary angles is 8 degrees less than the other.
Which of the following systems of equations represents the word problem?
x + y = 90 and x - y = 8
x + y = 90 and 90 - x = 8
x + y = 90 and y = 8x

Answer by checkley79(3341)

X+Y=90
X=Y-8 REPLACE X WITH (Y-8)
(Y-8)+Y=90
2Y=90-8
2Y=82
Y=82/2
Y=41 ANS.
X=41+8
X=49 ANS.
PROOF:
41+49=90
90=90